# Words as Trigger Points in Social Media Discussions

### v1

@@ Abstract
Trigger points are a concept introduced by Mau, Lux, and Westheuser (2023) to study qualitative focus group interviews and understand polarisation in Germany. When people communicate, trigger points represent moments when individuals feel that their understanding of what is fair, normal, or appropriate in society is questioned. In the original studies, individuals react affectively to such triggers and show strong and negative emotional responses. In this paper, we introduce the first systematic study of the large-scale effect of individual words as trigger points by analysing a large amount of social media posts. We examine online deliberations on Reddit between 2020 and 2022 and collect >>>100 million posts from subreddits related to a set of words identified as trigger points in UK politics. We find that such trigger words affect user engagement and have noticeable consequences on animosity in online discussions. We share empirical evidence of trigger words causing animosity, and how they provide incentives for hate speech, adversarial debates, and disagreements. Our work is the first to introduce trigger points to computational studies of online communication. Our findings are relevant to researchers interested in online harms and who examine how citizens debate politics and society in light of affective polarisation.
1 Introduction
@@
Trigger points, introduced by Mau, Lux, and Westheuser (2023), are rooted in theories of affective political identity and relate to deeply lying beliefs about moral expectations and social dispositions. Examining trigger points in online discussions helps understand why and when social media users engage in disagreements or affective political deliberations. This opens the door to modelling social media user engagement more effectively and studying the conditions and causal mechanisms that lead to negative reactions, hate speech, and abusive language in online debates.
@@
Figure 1: Consequences of trigger points. H1 refers to hypothesis 1, i.e., trigger words cause more engagement. H2 refers to hypothesis 2, i.e., trigger words cause more animosity.
@@
Trigger points are related to how individuals perceive themselves in society (Mau, Lux, and Westheuser 2023). People have subjective expectations about acceptable behaviour such as explicit knowledge of law-like rules and, more importantly, implicit understanding of appropriate actions. These deeply running beliefs of what is fair and unfair constitute a moral compass fundamental to navigating the social world. People feel triggered if these normative understandings are questioned. More formally, Mau, Lux, and Westheuser (2023, 246) define trigger points as “those moments that question what is acceptable in individuals understanding of the social contract. Trigger points jeopardise what individuals understand as the fabric of society and their own role in it.
Mau, Lux, and Westheuser (2023) propose four different mechanisms why trigger points affect individuals.
@@
Going against deeply held beliefs about individuals and their societal role causes discomfort. People respond to such triggers in an affective behavioural mode and increase a debate’s intensity and emotionality (Mau, Lux, and Westheuser 2023). In the social media context, we expect that this has two direct consequences (Figure 1). Trigger points lead to:
  • Hypothesis 1 more user engagement, which incites a higher number of messages;
  • Hypothesis 2 higher levels of animosity, which may lead to polarisation, negativity, anger, and hate.
In our study we focus on how individual words can already act as trigger points in online debates. We look into five trigger words (Rwanda, Brexit, NHS, Vaccine, Feminism) in the context of UK politics and collect >>>100 million related Reddit posts. We then leverage textual and Natural Language Processing (NLP) tools to look into the extent to which these words lead to potentially negative, emotional, and harmful responses, and analyse the results based on a systematic comparison between three treatment and control groups:
  1. 1.
    Space: specific subreddits we expect to be triggered by the selected words.
  2. 2.
@@
Our results show that these trigger words cause higher levels of engagement and more animosity, measured as an increase in the controversiality, negative sentiment, anger, and hate speech in a thread.
2 Related Work
Trigger points (Mau, Lux, and Westheuser 2023) bear great potential to understand polarisation and division in online discussions. However, the existence or validity of the concept has not been tested on a large scale. Trigger points can be particularly relevant to work on social media. They help understand the causes of disagreements in online conversations (De Kock and Vlachos 2021) and reasons for their escalation (De Kock and Vlachos 2022; van der Meer et al. 2023). Further to triggering ad hominem attacks, escalation can lead to hate speech. There is, however, no unanimous agreement on how it differs from abusive, and offensive language (Basile et al. 2019; Davidson et al. 2017; Talat and Hovy 2016).
Despite the rising interest in hate speech and harmful language in social media platforms, including its implicit forms (Ousidhoum et al. 2019; Sap et al. 2019; ElSherief et al. 2021; Ocampo et al. 2023), and dogwhistles, which are coded expressions that carry two meanings with a hateful or provocative coded one (Breitholtz and Cooper 2021), we know little about the sources of these phenomena. Examining what can trigger harmful reactions may help us make the detection process more robust and avoid the well-known problem of selection bias (Wiegand, Ruppenhofer, and Kleinbauer 2019; Ousidhoum, Song, and Yeung 2020) caused by the common data collection strategies based on hashtags (Talat and Hovy 2016) and keywords (Davidson et al. 2017).
Similarly, considering the broad array of harmful online behaviours, the wide range of elements they involve, and how they can be included in social media data (Olteanu et al. 2019), Blodgett et al. (2020) report a missing paradigm that helps assess the initial causes without mainly focusing on the performance of the detection model. Consequently, studying trigger points can help us track the origins of bias and assess whether a social media post can lead to hateful comments (Dahiya et al. 2021).
3 Methodology
Mau, Lux, and Westheuser (2023) developed the concept of trigger points based on qualitative interviews with focus groups to analyse the perceived polarisation of German society. We use this concept to study online communication on social media at a large scale. To critically assess its reliability, we test whether already individual words can trigger the reactions outlined in the two hypotheses, i.e., (1) more engagement and (2) animosity.
3.1 Term selection
@@
Table 1: Trigger words and their respective trigger mechanisms (described in the introduction).
Rwanda
@@
Rwanda plays a particular role in the UK’s immigration policy. [^1] https://www.gov.uk/government/publications/uk-rwanda-treaty-provision-of-an-asylum-partnership/uk-rwanda-treaty-provision-of-an-asylum-partnership-accessible Until late 2021, Rwanda was just the name of any other African country. However, with the then government, the term’s meaning changed in early 2022 and became emblematic of how the British government suggested handling asylum seekers, and the home secretary struck a deal with Rwanda: the Migration and Economic Development Partnership. Instead of handling asylum cases directly in the UK, migrants would be flown to Kigali, and their claims were to be processed there, in turn for U\$300m, (2.3% of Rwandan GDP) and expenses for every migrant sent (The Economist 2023). With this, Rwanda became a trigger word for discussions on immigration policy in the UK from 2022 onwards, particularly relating to concerns about “unfair favours towards migrants and unregulated immigration”.
Brexit
@@
The UK’s departure from the European Union has been controversial since its inception. Even eight years post-exit, debates persist regarding the soundness of the decision. The division among the population was already evident during the tightly contested 2016 referendum, which revealed distinct trends among various demographic groups based on location, age, and income (Arnorsson and Zoega 2018). Since the referendum, further divisions have emerged concerning the future relationship between the UK and EU (Vasilopoulou and Talving 2019). Additionally, misinformation and disinformation related to Brexit have intensified these debates (Höller 2021; Bruno, Lambiotte, and Saracco 2022). Here, considerations about treating EU migrants differently than the national population (inequality), norm transgressions of “criminal” migrants, particularly from the East of Europe, or opening the gate to “unlimited” immigration make Brexit a potential trigger word for analysis.
NHS
@@
Due to recent financial austerity measures, the UK government was accused of neglecting the health care system and reducing its budget, which leads to deteriorating services (The Guardian 2022; The Independent 2024). As a result, the NHS has become a hot topic of discussion among the UK population. Public opinion is divided, with some supporting that it receives adequate funding, while others believe it wastes resources or that its services are in decline (Gershlick, Charlesworth, and Taylor 2015). This trigger word is related to going against considerations about unequal treatment (i.e., inequality as shown in Table 1).
Vaccine
@@
Public vaccination programs, particularly those against COVID-19, have caused significant debate in recent years. Even after their widespread roll-out, discussions over their effectiveness and safety are ongoing. Initial reluctance was observed across various demographic segments, influenced by factors such as education level, political affiliation, and socioeconomic status (Troiano and Nardi 2021). Post-roll-out, debates have shifted to focus on issues such as the fairness of vaccine distribution and unreasonable demands for vaccination (Nichol and Mermin-Bunnell 2021). Moreover, the proliferation of misinformation and conspiracy theories regarding vaccines has further fuelled public division (Hayawi et al. 2022; Garett and Young 2021), particularly in social media, making Vaccine an ideal trigger word for our study.
@@
Table 2: Numbers, in millions, of threads, comments, and time period for each trigger word.
@@
Misogyny can be observed on social media, where anonymity often allows users to express offensive and hateful sentiments (Barker and Jurasz 2019). Previous work has shown that women politicians and journalists tend to be harassed online every few seconds (Posetti et al. 2022).  [^2] https://www.amnesty.org.uk/press-releases/women- abused-twitter-every-30-seconds-new-study Some social media influencers have built a large audience by sharing and promoting harmful content against women. This has led to the formation of online echo chambers (Farrell et al. 2019) where such views are amplified and propagated. Our research includes Feminism as a trigger word based on mechanisms related to norm transgressions, fears of a slippery slope and unreasonable behavioural demands.
@@
Table 3: Subreddits selected for each trigger word.
@@
Table 4: Percentage of comments per thread before and after each trigger appearance (Hypothesis 1). Difference in Difference (DiD) and 95% confidence interval (CI) included for the Control, Space and Semantics, comparisons. Bold scores indicate statistically significant results. Detailed distribution of comments can be seen in Appendix B, see Figure 7.
3.2 Data collection
Our study uses data from Reddit comments in English, collected via the Pushshift API (Baumgartner et al. 2020). We searched for exact matches of specific keywords in comments, ignoring the respective case, and gathered all corresponding posts from 2019 to 2022. Subsequently, the entire discussion thread for each comment extracted was collected [^3] When extracting comments for the term woman, we capped the maximum number of comments at 20 million due to the large volume.. We excluded posts in languages other than English, as identified with fastText-based language identifier (Bojanowski et al. 2017). Posts marked as “[deleted]” by the API were also removed. Following this, only threads that contained posts before and after the appearance of the target keywords were retained, as it is crucial to analyse the potential differences before and after the appearance of a given trigger point. As a final filtering step, and considering the fast-paced nature of Reddit, we limited our analysis to posts occurring 30 minutes before and 30 minutes after the first mention of the triggering word. Table 2 shows the total number of comments, threads, and comments with triggers collected for each trigger point and the time period they were gathered.
3.3 Research design
We identify the causal effect of trigger points with the difference-in-difference estimator (Angrist and Pischke 2009; Card and Krueger 1994). This estimator is based on systematically comparing a treatment and a control group. It calculates the causal effect of an intervention as the difference before and after the intervention in the treatment group minus the difference before and after the intervention in the control group. In our context, we expect users in the treatment group to be highly likely to be triggered by the respective terms and those in the control group to be much less triggered by the respective words.
Control groups
We systematically compare treatment and control groups along three different dimensions.
@@ 2. Semantics
We also compare trigger words with a term that shares a similar meaning without going against the deeper underlying transgressions that incite trigger responses. To members of the subreddits in Table 3, using an explicit trigger word or its more harmless control term should make a real difference. We define related control group words and contrast woman to Feminism/Feminist, Red Cross to NHS, EU/European Union to Brexit, surgery to vaccine, and names of Rwanda’s neighbouring countries [^4] Burundi, Tanzania, Congo, Democratic Republic of Congo, Uganda, Kenya, and Zambia. to Rwanda.
@@ 3. Time
In the third experiment, we make use of the fact that two of our terms—Vaccine and Rwanda—only became explicit trigger words at specific points in time. For “Rwanda” we compare discussions in subreddits likely to incite reactions in 2020 to discussions in 2022. For Vaccine we compare threads from 2019 (before the COVID-19 pandemic) to 2022’s (post COVID-19 pandemic). Again, we expect that users from the subreddits in Table 3 are triggered by those only in 2022 but not in the respective earlier years.
Statistical analysis
We use the difference-in-difference (DiD) estimator to analyse the causal effect of the trigger word on different online behaviours. The idea behind the DiD estimator is to compare the outcome change between the treatment and control groups before and after the trigger word. Specifically, we calculate the difference in the average outcome change for the treatment group before and after the trigger word and then subtract the difference in the average outcome change for the control group. If the treatment has a causal effect, we expect a larger change for the treatment group than for the control group. For each thread, we normalise by the actual number of messages, and calculate the share of messages before and after the appearance of the trigger word. To estimate the causal effect, we specify the DiD as an OLS regression (Angrist and Pischke 2009):

### v2

@@ Abstract.
Trigger points are a concept introduced by Mau et al. (2023) to study qualitative focus group interviews and understand polarisation in Germany. When people communicate, trigger points represent moments when individuals feel that their understanding of what is fair, normal, or appropriate in society is questioned. In the original studies, individuals show strong and negative emotional responses when certain triggering words or topics are mentioned. In this paper, we introduce the first systematic study of the large-scale effect of individual words as trigger points by analyzing a large amount of social media posts. We examine online deliberations on Reddit between 2020 and 2022 and collect >>>100 million posts from subreddits related to a set of words identified as trigger points in UK politics. We find that such trigger words affect user engagement and have noticeable consequences how online discussions occur. Trigger words cause animosity and incentivise hate speech, adversarial debates, and disagreements. Our work is the first to introduce trigger points to computational studies of online communication. The findings are relevant to researchers interested in affective computing, online deliberation, and those examining how citizens debate politics and society in light of affective polarisation.
Affective Computing, Polarization, Reddit, Trigger Words, User Engagement
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: THE WEB CONFERENCE 2025; 28 April–2 May, 2025; Sydney, Australiaccs: Information systems World Wide Webccs: Computing methodologies Natural language processing
1. Introduction
@@
Trigger points, introduced by Mau et al. (2023), are rooted in theories of affective political identity and relate to deeply lying beliefs about moral expectations and social dispositions. Examining trigger points in online discussions helps understand why and when social media users engage in disagreements or affective political deliberations. This opens the door to modelling social media user engagement more effectively and studying the conditions and causal mechanisms that lead to adverse reactions, hate speech, and abusive language in online debates.
@@
Figure 1. Consequences of trigger points. H1 refers to hypothesis 1, i.e., trigger words cause more engagement. H2 refers to hypothesis 2, i.e., trigger words cause more animosity.
@@
Trigger points are related to how individuals perceive themselves in society. People have subjective expectations about acceptable behaviour, such as explicit knowledge of law-like rules and, more importantly, an implicit understanding of appropriate actions. These deeply running beliefs of what is fair and unfair constitute a moral compass fundamental to navigating the social world. People feel triggered if these normative understandings are questioned. More formally, Mau et al. (2023, 246) define trigger points as “those moments that question what is acceptable against the backdrop of individuals’ understanding of the social contract. Trigger points jeopardise what individuals consider as the fabric of society and their role in it.” Trigger points can affect individuals through four mechanisms.
@@
Going against deeply held beliefs about individuals and their societal role causes discomfort. People respond to such triggers in an affective behavioural mode and increase a debate’s intensity and emotionality.
In the social media context, we expect that this has two direct consequences (Figure 1). Trigger points lead to:
  • Hypothesis 1 more user engagement, which incites a higher number of messages or comments;
  • Hypothesis 2 higher levels of animosity, which incite polarisation, negativity, anger, and hate.
Our study focuses on how individual words can act as trigger points in online debates. We look into five words (Rwanda, Brexit, NHS, Vaccine, Feminism) we expect to work as trigger points in the context of UK politics. Collecting >>>100 million related Reddit posts, we use text analytics and NLP methods to understand whether these words lead to potentially negative, emotional, and harmful responses. The analysis builds upon a systematic comparison between a treatment and control group in three different scenarios:
  1. (1)
    Space: specific subreddits where the selected words trigger users.
  2. (2)
@@
Our results show that the trigger words cause higher engagement and more animosity, measured as an increase in a thread’s controversiality, negative sentiment, anger, and hate speech.
2. Related Work
Trigger points (Mau et al., 2023) bear great potential to understand polarisation and division in online discussions. However, the existence or validity of the concept has not been tested on a large scale. Trigger points can be particularly relevant to work on social media. They help understand the causes of disagreements in online conversations (De Kock and Vlachos, 2021) and reasons for their escalation (De Kock and Vlachos, 2022; van der Meer et al., 2023). Further to triggering ad hominem attacks, escalation can lead to hate speech. There is, however, no unanimous agreement on how it differs from abusive and offensive language (Basile et al., 2019; Davidson et al., 2017; Talat and Hovy, 2016).
Despite the rising interest in hate speech and harmful language in social media platforms, including its implicit forms (Ousidhoum et al., 2019; Sap et al., 2019; ElSherief et al., 2021; Ocampo et al., 2023), and dog-whistles which are coded expressions that carry two meanings with a hateful or provocative coded one (Breitholtz and Cooper, 2021), we know little about the sources of these phenomena. Examining what can trigger harmful reactions may help us make more informed choices when collecting data, thus, leading to more robust detection models and mitigating the well-known problem of selection bias (Wiegand et al., 2019; Ousidhoum et al., 2020).
Similarly, considering the broad array of harmful online behaviours, the wide range of elements they involve, and how they can be included in social media data (Olteanu et al., 2019), Blodgett et al. (2020) report a missing paradigm that helps assess the initial causes without mainly focusing on the performance of the detection model. Studying trigger points helps track the origins of bias and assess whether a social media post can lead to hateful comments (Dahiya et al., 2021).
Further, understanding how trigger words affect online communication also sheds light on how language affects polarisation on social media (Karjus and Cuskley, 2024). It lays the groundwork for understanding how polarised rhetoric may emerge and lead to intergroup conflict (Simchon et al., 2022).
3. Methodology
We use the UK context as a large-scale case study to observe polarisation when individuals communicate on social media. Our goal is to report behavioural mechanisms similar to those observed by Mau et al. (2023) when conducting interviews to analyse the perceived polarisation of the German society. Based on a curated set of likely candidates, we test whether individual words can trigger the reactions outlined in the two hypotheses, i.e., (1) more engagement and (2) animosity.
3.1. Term selection
@@
Table 1. Trigger words and their respective trigger mechanisms (described in the introduction).
Rwanda.
@@
Rwanda plays a particular role in the UK’s immigration policy. [^1] https://www.gov.uk/government/publications/uk-rwanda-treaty-provision-of-an-asylum-partnership/uk-rwanda-treaty-provision-of-an-asylum-partnership-accessible Until late 2021, Rwanda was just the name of any other African country. However, in early 2022 the term’s meaning became emblematic of how the British government suggested handling asylum seekers, and the home secretary struck a deal with Rwanda: the Migration and Economic Development Partnership. Instead of processing asylum cases directly in the UK, migrants would be flown to Kigali, and their claims were to be assessed there, in turn for U\$300m, (2.3% of Rwandan GDP) and expenses for every migrant sent. With this, Rwanda became a trigger word for discussions on immigration policy in the UK from 2022 onwards, particularly related to “unfair favours towards migrants and unregulated immigration” (The Economist, 2023).
Brexit.
@@
The UK’s departure from the European Union has been controversial since its inception. Even eight years post-exit, debates persist regarding the soundness of the decision. The division among the population was already evident during the tightly contested 2016 referendum, which revealed distinct trends among various demographic groups based on location, age, and income (Arnorsson and Zoega, 2018). Since the referendum, further divisions have emerged concerning the future relationship between the UK and EU (Vasilopoulou and Talving, 2019). Additionally, misinformation and disinformation have intensified these debates (Höller, 2021; Bruno et al., 2022). Here, considerations about treating EU migrants differently than the national population (inequality), norm transgressions of “criminal” migrants, particularly from the East of Europe, or opening the gate to “unlimited” immigration make Brexit a potential trigger word for analysis.
NHS.
@@
Due to financial austerity measures, the UK government was accused of reducing its budget and neglecting the health care system (The Guardian, 2022; The Independent, 2024). As a result, the NHS has become a hot topic of discussion among the UK population. Public opinion is divided, with some supporting that it receives adequate funding, while others believe it wastes resources or that its services are in decline (Gershlick et al., 2015). This trigger word is related to going against considerations about unequal treatment (i.e., inequality as shown in Table 1).
Vaccine.
@@
Public vaccination programs, particularly those against COVID-19, have caused significant debate in recent years. Even after their widespread roll-out, discussions over their effectiveness and safety are ongoing. Initial reluctance was observed across various demographic segments, influenced by factors such as education level, political affiliation, and socioeconomic status (Troiano and Nardi, 2021). Post-roll-out, debates have shifted to focus on issues such as the fairness of vaccine distribution and unreasonable demands for vaccination (Nichol and Mermin-Bunnell, 2021). Moreover, the proliferation of misinformation and conspiracy theories has further fuelled public division (Hayawi et al., 2022; Garett and Young, 2021), particularly in social media, making Vaccine an ideal trigger word for our study.
@@
Table 2. Numbers, in millions, of threads, comments, and time period for each trigger word.
@@
Misogyny can be observed on social media, where anonymity often allows users to express offensive and hateful sentiments (Barker and Jurasz, 2019). Previous work has shown that women politicians and journalists tend to be harassed online every few seconds (Posetti et al., 2022).  [^2] https://www.amnesty.org.uk/press-releases/women- abused-twitter-every-30-seconds-new-study Some social media influencers have built a large audience by sharing and promoting harmful content against women. This has led to the formation of online echo chambers (Farrell et al., 2019) where such views are amplified and propagated. Our research includes Feminism as a trigger word based on mechanisms related to norm transgressions, fears of a slippery slope and unreasonable behavioural demands.
@@
Table 3. Subreddits selected for each trigger word.
@@
Table 4. Percentage of comments per thread before and after each trigger appearance (Hypothesis 1). Difference-in-Difference (DiD) and 95% confidence interval (CI) included for the Control, Space and Semantics, comparisons. Bold scores indicate statistically significant results. Detailed distribution of comments can be seen in Appendix B, see Figure 7.
3.2. Data collection
Our study uses data from Reddit comments in English, collected via the Pushshift API (Baumgartner et al., 2020). We searched for exact matches of specific keywords in comments from 2019 to 2022, and then collected the entire discussion thread for each extracted comment [^3] When extracting comments for the term woman, we capped the maximum number of comments at 20 million due to the large volume.. We excluded posts in languages other than English, as identified with fastText-based language identifier (Bojanowski et al., 2017). Posts marked as “[deleted]” by the API were also removed. Following this, only threads that contained posts before and after the appearance of the target keywords were retained, as it is crucial to analyse the potential differences before and after the appearance of a given trigger point.
As a final filtering step we limited our analysis to posts occurring 30 minutes before and 30 minutes after the first mention of the triggering word. Previous research has shown that Reddit posts typically receive the majority of interactions within a short window after the post is created (Thukral et al., 2018). Moreover, early responses are often important in shaping the direction of the conversation (Singer et al., 2016; Horne et al., 2017). Considering the fast-paced nature of Reddit, the 30 minutes window filters out unrelated responses and ensures the focus of our analysis remains on the most relevant data. Table 2 shows the total number of comments, threads, and comments collected for each trigger word and the period they were gathered.
3.3. Research design
We identify the causal effect of trigger points with the difference-in-difference estimator (DiD) which systematically compares a treatment with a control group (Angrist and Pischke, 2009; Card and Krueger, 1994). It calculates the causal effect of an intervention as the difference before and after the intervention in the treatment group minus the difference before and after the intervention in the control group. In our context, we expect users in the treatment group to be highly likely to be triggered by the respective terms and those in the control group to be much less triggered by the respective words. There should be a more substantive behavioural change for the treatment group than for the control group.
3.3.1. Control groups
We systematically compare treatment and control groups along three different dimensions. [^4] Identifying the causal effect in three completely unrelated comparisons helps address concerns about eventual endogeneity.
@@ 2. Semantics
We also compare trigger words with a term that shares a similar meaning, but that does not go against the deeper underlying transgressions responsible for trigger responses. To members of the subreddits in Table 3, using an explicit trigger word or its more harmless control term should make a real difference. We define related control group words and contrast woman to Feminism/Feminist, Red Cross to NHS, EU/European Union to Brexit, surgery to vaccine, and names of Rwanda’s neighbouring countries [^5] Burundi, Tanzania, Congo, Democratic Republic of Congo, Uganda, Kenya, and Zambia. to Rwanda.
@@ 3. Time
In the third experiment, we use the fact that two of our terms—Vaccine and Rwanda—only became explicit trigger words at specific points in time. For “Rwanda” we compare discussions in subreddits likely to incite reactions in 2020 to discussions in 2022. For Vaccine we compare threads from 2019 (before the COVID-19 pandemic) to 2022’s (post COVID-19 pandemic). Again, we expect that users from the subreddits in Table 3 are triggered by those only in 2022 but not in the respective earlier years.
3.3.2. Statistical analysis
We estimate the trigger words’ treatment effect in 60 separate experiments—12 experiments for H1 (Table 4) and 48 experiments for H2 (Table 5). For each experiment, we prepare the data in the same way. The treatment/placebo is the first mention of a trigger word in a discussion thread, i.e., there is only one well-defined treatment per thread. To test H1, we count the number of posts for each discussion thread 30 minutes before and 30 minutes after the treatment in the treatment group and the placebo treatment in the control group. We then calculate the share of posts before and after the intervention. For example, if there are 8 posts during the 30 minutes before the trigger word and 12 posts during the 30 minutes after the trigger word, we normalise the data of this discussion thread to 40% before intervention and 60% after intervention. For experiments with H2, we count the number of the respective animosity in a thread and then normalise the data as a share of animosity per thread before and after the trigger word.
We estimate the causal effect in an OLS-regression framework. More specifically, we specify the DiD with the two-way fixed effects estimator (Angrist and Pischke, 2009):